OpenStudy (anonymous):

A tea shop owner is mixing a blend of two teas, one of which costs $6.50 per pound, the other costing $4.00 per pound. The owner wants to have 20 pounds of a mixture that will sell for $5.50 per pound. How much of each type of tea should be used?

OpenStudy (anonymous):

let x = pounds of 6.50 tea and let 20 - x = pounds of 4 tea 6.50x + 4(20-x) = 110 6.50x+ 80 - 4x = 110 -80 -80 2.50x = 30 /2.50 x=12 so, 12 pounds of $6.50 and 8 pounds of $4 tea.